OpenStudy (anonymous):

HELP PLEASE! If the maximum kinetic energy of a 1 kg mass on a spring with a spring constant of 120 N/m is 12.5 J with a period of 6 seconds, what is the amplitude of this motion?

OpenStudy (anonymous):

\[1/2mv _{\max}^{2}=1/kx _{\max}^2=1/2kA^{2}\]Solve for A

OpenStudy (anonymous):

Don't see it? All energy is kinetic when K is max, all energy is in the spring when A is max so they are equal to each other (conservation of energy)